OpenStudy (anonymous):

Find y. y'-e^ysinx=0

OpenStudy (bahrom7893):

another diff eqn?

OpenStudy (anonymous):

yess

OpenStudy (bahrom7893):

yay lol workin on it.

OpenStudy (bahrom7893):

FIRST OF ALL, In ALL Differential equations, especially if they are in terms of x and y, not only one variable, rewrite y' as dy/dx

OpenStudy (bahrom7893):

dy/dx - (e^y)(Sinx) = 0 dy/dx = (e^y)(Sinx)

OpenStudy (bahrom7893):

Divide both sides by e^y and multiply by dx: dy/e^y = Sinx dx

OpenStudy (bahrom7893):

For first integral: rewrite dy/e^y as e^(-y)dy, then let u = -y; du = -dy

OpenStudy (bahrom7893):

So: Integral of (e^(-y)dy) = - Integral of (e^(-y)(-dy)) [I multiplied by two minuses ( double negative is a positive) to get a -dy=du]

OpenStudy (bahrom7893):

- Integral of (e^(-y)(-dy)) = - Integral of (e^u du) = - e^u = - e^(-y)

OpenStudy (bahrom7893):

Int (dy/e^y)= Int (Sinx dx) - e^(-y) = - Cos(x) + C

OpenStudy (bahrom7893):

multiply everything by -1 ( - 1 times a constant is still a constant so I will let -C be A) e^(-y) = Cos(x) + A

OpenStudy (bahrom7893):

Take Ln of both sides: Ln(e^(-y)) = Ln(Cos(x)+A) -y = Ln(Cos(x)+A) y = - Ln(Cos(x)+A)